1.Define the diffusion potential. What is a salt bridge and what is it used for?

b-) One of the concentration cell was given as Ni I Ni+21 II Ni+22 II Ni with salt bridge, wherein the activities a Ni2+,I = 0. 7 and a Ni2+,II = 1. T = 25°C. Calculate the EMF of the cell if the electrochemical potentials are the same.

2.How large should the ratio of the activities of the participating metal ions (which are they) be in a Daniell - Cell at 298 K for the characteristic reaction of the Daniell element to occur spontaneously? (please use also necessary equations)
Answer #1

1. a) Diffusion potential is the potential difference generated across a membrane because of a concentration difference of an ion. It can be generated only if the membranes is permeable to the ion. The size of diffusion potential depends on the size of concentration gradient.

A salt bridge is used to connect the internal circuit of the oxidation and reduction electrodes of a galvanic cell. It's used to maintain electrical neutrality.

b) Since electrochemical potentials are the same, E​​​​​​0​​​​​ = 0V.

E = -RTlnK

K = (activity of Ni​​​​​2+)/(activity of Ni)

= 0.7/1

=0.7

E = -8.314*298ln0.7

=883.68 V

2. Take E = 1.1018 V ( voltage of daniel cell )

R= 8.314 J/mol

T= 298K

=> lnK = - 1.1018/(8.314*298)

= -0.0004417

=> K = 0.9995 ~ 1.

Thus since both are present in equimolar proportions, the ratio of activities of the participating ions is also equal to 1.